OpenStudy (anonymous):

if set A={1,1,2,3} B={1,1,5,7} is AintersectionB={1,1} or just {1} ? is AUB={1,2,3,5,7}

jimthompson5910 (jim_thompson5910):

you would remove all duplicates, so A intersect B = {1}

jimthompson5910 (jim_thompson5910):

and AUB = {1,2,3,5,7}, so you are correct there

OpenStudy (anonymous):

oh okay so you remove all duplicates what if it was like probability? e.g. dice roll does it stay the same

OpenStudy (anonymous):

can you have 2 same elemnets in a set

jimthompson5910 (jim_thompson5910):

what do you mean can you elaborate

OpenStudy (anonymous):

if A person rolled 1,1,4,5 and B rolled 1,1,5,10 does A intersectB represent anything??

OpenStudy (anonymous):

oh and what if you wanna put it on venn diagram

OpenStudy (anonymous):

do you only put a 1 in intersection or 2

jimthompson5910 (jim_thompson5910):

A intersect B would just be the set {1} and that would tell us that A and B both rolled 1 at some point

OpenStudy (anonymous):

oh okay veen diagram?

jimthompson5910 (jim_thompson5910):

oh wait A = {1,1,4,5} B = {1,1,5,10} A intersect B = {1,5}

jimthompson5910 (jim_thompson5910):

would look like this 

jimthompson5910 (jim_thompson5910):

but you're removing duplicates, so it becomes this 

OpenStudy (zzr0ck3r):

we do not list an element in a set more than once.

OpenStudy (anonymous):

A={1,1,4,5} , represent what A person rolled on a dice Pr(1)=1/2?? or 1/3

OpenStudy (anonymous):

or cant be {1,1,4,5}

jimthompson5910 (jim_thompson5910):

There are two '1' rolls out of 4, so P(1) = 2/4 = 1/2 for person A the same can be said for person B this is probably why it's a bad idea to use sets in this case

OpenStudy (anonymous):

ah okay but so can you list element twices?

jimthompson5910 (jim_thompson5910):

no each element is listed once in a set

OpenStudy (anonymous):

no{1,1,4,6} yes{1,4,6}?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

but then how you do probability if the numbers represent a dice roll